# On Bayesian Neural Networks with Dependent and Possibly Heavy-Tailed Weights

### v4

@@ Abstract
We consider fully connected and feedforward deep neural networks with dependent and possibly heavy-tailed weights, as introduced in [26], to address limitations of the standard Gaussian prior. It has been proved in [26] that, as the number of nodes in the hidden layers grows large, according to a sequential and ordered limit, the law of the output converges weakly to a Gaussian mixture. Among our results, we present sufficient conditions on the model parameters (the activation function and the associated Lévy measures) which ensure that the sequential limit is independent of the order. Next, we study the neural network through the lens of the posterior distribution with a Gaussian likelihood. If the random covariance matrix of the infinite-width limit is positive definite under the prior, we identify the posterior distribution of the output in the wide-width limit according to a sequential regime. Remarkably, we provide mild sufficient conditions to ensure the aforementioned invertibility of the random covariance matrix under the prior, thereby extending the results in [8]. We illustrate our findings using numerical simulations.
MSC 2010 Subject Classification: 68T07; 68W20; 60B10.
@@
Keywords: Deep Neural Networks, Infinite-Width Limit, Gaussian Mixture, Random Matrices.
1 Introduction
@@
A fully connected and feedforward neural network (which we simply refer to as a “neural network”  throughout the paper) is the simplest neural network architecture and consists of a sequence of hidden layers stacked between an input layer and an output layer. Each node (or neuron) in a layer is connected to all the nodes in the subsequent layer. These networks are used to estimate unknown functions relating observed inputs to outputs. Once the parameters of the network, i.e. , biases and weights, have been estimated on the basis of a training dataset, network’s output is a good approximation of the unknown target function. A neural network is called “deep”  if it has more than one hidden layer and “shallow”  if it has only one (we refer the reader to [17] for an introduction to neural networks).
@@
The Bayesian approach to the analysis of neural networks allows to include in the model both a prior knowledge on the parameters and the observed data through a prior distribution on neural network’s parameters and a likelihood function, respectively. Neal [32, 33] initiated the theoretical study of Bayesian neural networks by proving that, if a Bayesian shallow neural network is initialized with independent Gaussian parameters (i.e. , the prior is Gaussian), then the output of the neural network converges in distribution to a Gaussian process as the number of neurons in the hidden layers increases, i.e. , in the infinite-width limit. This result was extended to Bayesian deep neural networks two decades later (see [19, 25, 30]), and has only recently been made quantitative using optimal transport theory (see [6, 37]), the Stein method for Gaussian approximation (see [3, 4, 14]), and alternative techniques ([7, 11]). Another promising approach to analyze Bayesian neural networks is through the lens of large deviations. First results in this direction are given in [20, 27]. A different perspective is provided by the so-called mean field analysis of neural networks (see [18, 31]).
@@
The emergence of Gaussian processes improved our understanding of how large neural networks work and how to make them more efficient. It also motivated the use of Bayesian regression inference methods, see [25]. However, as noticed in [32] and [26], the connection with Gaussian processes also highlighted the limitations of Bayesian neural networks with a Gaussian prior. Indeed, there are at least three drawbacks with the choice of independent Gaussian weights: $(i)$ Hidden layers do not represent hidden features that capture important aspects of the data; $(ii)$ In the infinite-width limit the coordinates of the output become independent and identically distributed Gaussian processes, which is usually undesirable; $(iii)$ The assumption of independent Gaussian weights is often unrealistic, as estimated weights of deep neural networks show dependencies and heavy tails (see [16, 29, 38]).
@@
To circumvent these limitations, some authors ([5, 13, 23]) proposed priors which account for independent and identically distributed non-Gaussian weights. However, due to the independence assumption, in the infinite-width limit, the output of the neural network still converges to a stochastic process with independent coordinates. A more structured prior on the weights has been proposed by [1, 26]. Letting $W_{hj}^{(\ell+1)}$ denote the random weight between the node $h$ at the layer $\ell+1$ and the node $j$ at the layer $\ell$, in [26] it is assumed that $W_{hj}^{(\ell+1)}:=\sqrt{V_{n_{\ell},j}^{(\ell)}}N_{hj}^{(\ell)}$, where $V_{n_{\ell},j}^{(\ell)}$ are independent random variances, which are identically distributed over $j$, and $N_{hj}^{(\ell)}$ are independent and Gaussian distributed random variables, with mean 0 and variance $C_{W}>0$, independent of the random variances. We refer the reader to Section 2.2 for a rigorous description of the model. The prior proposed in [26] is more general than the Gaussian one, which is retrieved setting $V_{n_{\ell},j}^{(\ell)}:=1/n_{\ell}$, and it accounts for dependent and heavy-tailed weights. Indeed, for fixed $\ell$ and $j$, the weights $\{W_{hj}^{(\ell+1)}\}_{h}$ are stochastically dependent and, if the random variances are distributed according to the square of a heavy-tailed distribution with support on $(0,\infty)$, a simple computation shows that the corresponding weights are also heavy-tailed. It is proved in [26] that if the aggregate random variances at the level of the layer $\ell+1$ converge in distribution to an infinitely divisible law, as the number of nodes in the layer $\ell$ grows large, then the output of the neural network converges in law to a mixture of Gaussian processes whose coordinates are dependent but still identically distributed. Such a convergence takes place as the number of nodes in the hidden layers grows large, according to a sequential limit with a prescribed order (see Theorem 16 in [26] or Theorem 4.1 for the precise statement). In [26] it is left as challenging open problem to find more natural “limiting schemes”  for the validity of such a convergence to a Gaussian mixture. Theorem 4.2 (whose proof is in Appendix) establishes sufficient conditions for the sequential limit of Theorem 16 in [26] to be independent of the order. This generalization is achieved via an alternate representation of the neural network (see Lemma 7.1), which has been used in [27] to study the large deviations of the output in the case of a Gaussian prior.
@@
In this paper we are concerned with posterior Bayesian neural networks with dependent and possibly heavy-tailed weights $\{W_{hj}^{(\ell+1)}\}$ as described above. If the likelihood is Gaussian, we identify the infinite-width posterior distribution of the output and show that, under a suitable probability measure, it is a Gaussian mixture whose coordinates are dependent and not identically distributed (see Theorem 5.1). This result entirely circumvents one of the limitations arising from the choice of a Gaussian prior. Such a result holds, however, under the assumption that the random covariance matrix of the infinite-width limit of the output is invertible under the prior distribution. Remarkably, we provide sufficient conditions that guarantee such an invertibility assumption, extending the results in [8] to a more general framework. (see Theorems 6.1 and 6.2). The posterior law of the output of the Bayesian neural network at a finite width is very complicated, especially for large values of the depth. One must resort to simulations to gain insight into this distribution. In the final section of the paper, we provide numerical illustrations of our main result.
@@ 2 Neural networks
2.1 Neural networks and statistical learning
From a mathematical point of view, fully connected and feedforward neural networks are a parametrized family of functions, recursively defined as follows. Let $L$, $n_{0}$,…, $n_{L+1}\in\mathbb{N}^{*}$ be integers. For $\ell=1,\ldots,L$, we set
$Z_{h}^{(\ell+1)}(x)$ $:=B_{h}^{(\ell+1)}+\sum_{j=1}^{n_{\ell}}W_{hj}^{(\ell+1)}\sigma(Z_{j}^{(\ell)}(x)),\quad\text{$h=1,\ldots,n_{\ell+1}$}$
$Z_{h}^{(1)}(x)$ $:=B_{h}^{(1)}+\sum_{j=1}^{n_{0}}W_{hj}^{(1)}x_{j},\quad\text{$h=1,\ldots,n_{1}$}$
where $x\in\mathbb{R}^{n_{0}}$ is the input, $\sigma:\mathbb{R}\to\mathbb{R}$ (a measurable function) is the activation function, $\{B_{h}^{(\ell)}\}_{\ell=1,\ldots,L+1;h=1,\ldots,n_{\ell}}$ and $\{W_{hj}^{(\ell)}\}_{\ell=1,\ldots,L+1;h=1,\ldots,n_{\ell};j=1,\ldots,n_{\ell-1}}$ are real parameters called biases and weights, respectively. The network consists of $L$ hidden layers stacked between the input layer (the layer 0) and the output layer (the layer $L+1$). On each hidden layer $\ell\in\{1,\ldots,L\}$, there are $n_{\ell}$ nodes. The network is called deep if $L\geq 2$ and shallow if $L=1$. Throughout the paper, we often apply $\sigma$ to vectors, meaning that $\sigma(x)=(\sigma(x_{1}),\ldots,\sigma(x_{p}))^{\top}$ for $x=(x_{1},\ldots,x_{p})^{\top}$.
In statistical learning, neural networks are used to estimate unknown target functions. More precisely, for a fixed network architecture ($L$, $n_{0},\ldots,n_{L+1}$, $\sigma$), for a given unknown target function $f:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{L+1}}$ and a training dataset $\mathcal{D}:=\{(x(i),y(i))\}_{i=1,\ldots,d}\subset\mathbb{R}^{n_{0}}\times\mathbb{R}^{n_{L+1}}$, $d\in\mathbb{N}^{*}$, (i.e., couples of data $x(i)$ and outcomes $y(i):=f(x(i))$) the objective is to produce an estimate of the parameter $\Theta=(\{B_{h}^{(\ell)}\},\{W_{hj}^{(\ell)}\})$, say $\widehat{\Theta}$, in such a way that the output of the neural network is a good estimate of $f$, i.e.,
@@ 2.1 Neural networks and statistical learning
where
$\mathcal{T}:=\{(x^{\prime}(j),y^{\prime}(j))\}_{j=1,\ldots,t}\subset\mathbb{R}^{n_{0}}\times\mathbb{R}^{n_{L+1}}$, $t\in\mathbb{N}^{*}$
is a test set. Throughout the paper we encode the inputs (data) in the $n_{0}\times d$ matrix $\mathbb{x}:=(x(1)\ldots x(d))$ and the outcomes (responses) in the $n_{L+1}\times d$ matrix $\mathbb{y}:=(y(1)\ldots y(d))$.
2.2 Neural networks with dependent and possibly heavy-tailed weights, and Bayesian statistical learning
@@
  • For $\ell=1,\ldots,L+1$ and $h=1,\ldots,n_{\ell}$, $B_{h}^{(\ell)}$ are random variables with $B_{h}^{(\ell)}\sim\mathcal{N}_{1}(0,C_{B}),$ for a constant $C_{B}\geq 0$;
  • For $\ell=1,\ldots,L+1$, $h=1,\ldots,n_{\ell}$ and $j=1,\ldots,n_{\ell-1}$, $W_{hj}^{(\ell)}$ are random variables defined by
    $W_{hj}^{(\ell)}:=\sqrt{V_{n_{\ell-1},j}^{(\ell-1)}}N_{hj}^{(\ell)},$
    where: for $j=1,\ldots,n_{0}$, $V_{n_{0},j}^{(0)}:=n_{0}^{-1}$, for $\ell=2,\ldots,L+1$, $\{V_{n_{\ell-1},j}^{(\ell-1)}\}_{j=1,\ldots,n_{\ell-1}}$ are non-negative and identically distributed (over $j$) random variables, and for $\ell=1,\ldots,L+1$, $h=1,\ldots,n_{\ell}$ and $j=1,\ldots,n_{\ell-1}$, $N_{hj}^{(\ell)}$ are random variables with $N_{hj}^{(\ell)}\sim\mathcal{N}_{1}(0,C_{W}),$ for a constant $C_{W}>0$;
  • All the random variables $\{B_{h}^{(\ell)},V_{n_{\ell-1},j}^{(\ell-1)},N_{hj}^{(\ell)}\}$ are stochastically independent.
Under these distributional assumptions, one speaks of neural network with dependent and possibly heavy-tailed weights, see the seminal paper [26], where this prior has been introduced. Indeed, note that, for a fixed $\ell\in\{2,\ldots,L+1\}$ and $j\in\{1,\ldots,n_{\ell-1}\}$, the random weights $W_{1j}^{(\ell)},\ldots,W_{n_{\ell}j}^{(\ell)}$ are dependent, and that, if $V_{n_{\ell-1},j}^{(\ell-1)}$ is distributed as the square of a random variable with a heavy tail law with support on $(0,\infty)$, then the distribution of $W_{hj}^{(\ell)}$ is heavy tail. Note also that if $V_{n_{\ell-1},j}^{(\ell-1)}:=\frac{1}{n_{\ell-1}}$, then we recover the usual Gaussian prior.
The Bayesian appraoch to the statistical learning problem allows to incorporate in the model the training dataset $\mathcal{D}$ through a likelihood function $\mathcal{L}(\mathcal{D},\Theta)$. Then the posterior knowledge on the parameter is summarized by the posterior probability measure
@@
where $\mathbb{E}_{\rm prior}$ denotes the expectation under the prior probability measure and it is assumed $\mathbb{E}_{\rm prior}\mathcal{L}(\mathcal{D},\Theta)>0.$ If the law of $\Theta$ under the prior has density $\mathrm{p}_{prior}(\cdot)$, then the law of $\Theta$ under the posterior has density
$\mathrm{p}_{{\rm posterior}}(\theta)\propto\mathcal{L}(\mathcal{D},\theta)\mathrm{p}_{{\rm prior}}(\theta),$
and one estimates the parameter $\Theta$ by maximizing $\mathrm{p}_{{\rm posterior}}(\cdot)$ via an adequate variant of the Stochastic Gradient Descent (maximum a posteriori estimate), see e.g. [16].
3 Preliminaries

### v5

# Posterior Bayesian Neural Networks with Dependent Weights

@@ Abstract.
We consider fully connected and feedforward deep neural networks with dependent and possibly heavy-tailed weights, as introduced in [26], to address limitations of the standard Gaussian prior. It has been proved in [26] that, as the number of nodes in the hidden layers grows large, according to a sequential and ordered limit, the law of the output converges weakly to a Gaussian mixture. In this paper, we study the neural network through the lens of the posterior distribution with a Gaussian likelihood. If the random covariance matrix of the infinite-width limit is positive definite under the prior, we identify the posterior distribution of the output in the wide-width limit according to a sequential regime. Remarkably, we provide mild sufficient conditions to ensure the aforementioned invertibility of the random covariance matrix under the prior, thereby extending the results in [8]. Among our results, we present sufficient conditions on some model parameters (the activation function and the associated Lévy measures) which ensure that the sequential limits are independent of the order. We illustrate our findings with examples and numerical simulations.
@@
Keywords Bayesian Analysis; Deep Neural Networks; Gaussian Mixture; Infinitely Divisible Distributions; Random Matrices
2020 Mathematics Subject Classification 60F05; 62F15,68T07
1. Introduction
@@
A fully connected and feedforward neural network (which we simply refer to as a “neural network”  throughout the paper) is the simplest neural network architecture and consists of a sequence of hidden layers stacked between an input layer and an output layer. Each node (or neuron) in a layer is connected to all the nodes in the subsequent layer. These networks are used to estimate unknown functions relating observed inputs to outputs. Once the parameters of the network, i.e., biases and weights, have been estimated on the basis of a training dataset, the network itself is a good approximation of the unknown target function. We refer the reader to [17] for an introduction to this topic.
Overview.
@@
The Bayesian approach to the analysis of neural networks allows to include in the model both a prior knowledge on the parameters and the observed data through a prior distribution on neural network’s parameters and a likelihood function, respectively. Neal [32, 33] initiated the theoretical study of Bayesian neural networks by proving that, if a Bayesian shallow neural network is initialized with independent Gaussian parameters (i.e., the prior is Gaussian), then the output of the neural network converges in distribution to a Gaussian process as the number of neurons in the hidden layer increases, namely, in the infinite-width limit. This result was extended to deep neural networks two decades later [19, 25, 30] and has only recently been made quantitative using optimal transport theory [6, 36], the Stein method for Gaussian approximation [3, 4, 14], and alternative techniques [7, 11]. Another promising approach to analyze Bayesian neural networks is through the lens of large deviations. First results in this direction are given in [20, 27]. A different perspective is provided by the so-called mean field analysis of neural networks, see [18, 31].
@@
The emergence of Gaussian processes improved our understanding of how large neural networks work and how to make them more efficient. It also motivated the use of Bayesian regression inference methods, see [25]. However, as noticed in [32] and [26], the connection with Gaussian processes also highlighted the limitations of Bayesian neural networks with a Gaussian prior. Indeed, there are at least three drawbacks with the choice of independent Gaussian weights: $(i)$ Hidden layers do not represent hidden features that capture important aspects of the data; $(ii)$ In the infinite-width limit the coordinates of the output become independent and identically distributed Gaussian processes, which is usually undesirable; $(iii)$ The assumption of independent Gaussian weights is often unrealistic, as estimated weights of deep neural networks show dependencies and heavy tail, see [16, 29, 37].
@@
To circumvent these limitations, some authors [5, 13, 23] proposed priors which account for independent and identically distributed (i.i.d. for short) non-Gaussian weights. However, due to the independence assumption, in the infinite-width limit, the output of the neural network still converges to a stochastic process with independent coordinates. A more structured prior on the weights has been proposed by [1, 26]. Letting $W_{hj}^{(\ell+1)}$ denote the random weight between the node $h$ at the layer $\ell+1$ and the node $j$ at the layer $\ell$, in [26] it is assumed that $W_{hj}^{(\ell+1)}:=\sqrt{V_{n_{\ell},j}^{(\ell)}}N_{hj}^{(\ell)}$, where $V_{n_{\ell},j}^{(\ell)}$ are independent random variances, which are identically distributed over $j$, and $N_{hj}^{(\ell)}$ are independent and Gaussian distributed random variables (r.v.’s for short), with mean 0 and variance $C_{W}>0$, independent of the random variances. We refer the reader to Section 2.2 for a rigorous description of the model. The prior proposed in [26] is more general than the Gaussian one, which is retrieved setting $V_{n_{\ell},j}^{(\ell)}$ $1/n_{\ell}$, and it accounts for dependent and heavy-tailed weights. Indeed, for fixed $\ell$ and $j$, the weights $\{W_{hj}^{(\ell+1)}\}_{h}$ are stochastically dependent and, if the random variances are distributed according to the square of a heavy-tailed distribution with support on $(0,\infty)$, a simple computation shows that the corresponding weights are also heavy-tailed.
It is proved in [26] that if the aggregate random variances at the level of the layer $\ell+1$ converge in distribution (necessarily to an infinitely divisible law), as the number of nodes in the layer $\ell$ grows large, then the output of the neural network converges in law to a mixture of Gaussian processes whose coordinates are dependent but still identically distributed. Such a convergence takes place as the number of nodes in the hidden layers grows large, according to a sequential limit with a prescribed order (see Theorem 16 in [26] or Theorem 1 for the precise statement). In [26] it is left as challenging open problem to find more natural “limiting schemes”  for the validity of such a convergence to a Gaussian mixture.
@@
Our contribution.
In this paper we are concerned with Bayesian neural networks with dependent and possibly heavy-tailed weights $\{W_{hj}^{(\ell+1)}\}$ as described above.
In Theorem 2, we establish sufficient conditions for the sequential limit of [26, Theorem 16] to be independent of the order. As detailed in Appendix A, this generalization can be achieved via an alternate representation of the neural network, which has been used in [27] to study the large deviations of the output in the case of a Gaussian prior.
As our main result, in Theorem 3, for Gaussian likelihoods we identify the infinite-width posterior distribution of the output and we show that, under a suitable probability measure, it is a Gaussian mixture whose coordinates are dependent and not identically distributed. This result entirely circumvents one of the limitations arising from the choice of a Gaussian prior. Such a result holds, however, under the assumption that the random covariance matrix of the infinite-width limit of the output is invertible under the prior distribution.
Remarkably, we provide sufficient conditions that guarantee such an invertibility assumption, by extending the results in [8] to a more general framework, see Theorems 4 and 5. These results are achieved by combining geometric measure-theoretic tools and an elementary, but tricky, linear algebraic construction.
The posterior law of the output of the Bayesian neural network at a finite width is very complicated, especially for large values of the depth. One must resort to simulations to gain insight into this distribution. In the final section of the paper, we provide some numerical illustrations of our main result.
Plan of the paper
@@ 2. Neural networks
2.1. Neural networks and statistical learning
From a mathematical point of view, fully connected and feedforward neural networks are a parametrized family of functions, recursively defined as follows. Let $L$, $n_{0}$,…, $n_{L+1}\in\mathbb{N}^{*}$ be integers. For $\ell\in[L]$, we set
$Z_{h}^{(\ell+1)}(x)$ $:=B_{h}^{(\ell+1)}+\sum_{j=1}^{n_{\ell}}W_{hj}^{(\ell+1)}\sigma(Z_{j}^{(\ell)}(x)),\quad\text{$h\in[n_{\ell+1}]$}$
$Z_{h}^{(1)}(x)$ $:=B_{h}^{(1)}+\sum_{j=1}^{n_{0}}W_{hj}^{(1)}x_{j},\quad\text{$h\in[n_{1}]$}$
where $x\in\mathbb{R}^{n_{0}}$ is the input, $\sigma:\mathbb{R}\to\mathbb{R}$ (a measurable function) is the activation function, $\{B_{h}^{(\ell)}\}_{\ell\in[L+1];h\in[n_{\ell}]}$ and $\{W_{hj}^{(\ell)}\}_{\ell\in[L+1];h\in[n_{\ell}];j\in[n_{\ell-1}]}$ are real parameters called biases and weights, respectively. The network consists of $L$ hidden layers stacked between the input layer (the layer 0) and the output layer (the layer $L+1$). On each hidden layer $\ell\in[L]$, there are $n_{\ell}$ nodes. The network is called deep if $L\geq 2$ and shallow if $L=1$. Throughout the paper, we often apply $\sigma$ to vectors, meaning that $\sigma(x)=(\sigma(x_{1}),\ldots,\sigma(x_{p}))^{\top}$ for $x=(x_{1},\ldots,x_{p})^{\top}$.
In statistical learning, neural networks are used to estimate unknown target functions. More precisely, for a fixed network architecture ($L$, $n_{0},\ldots,n_{L+1}$, $\sigma$), for a given unknown target function $f:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{L+1}}$ and a training dataset $\mathcal{D}:=\{(x(i),y(i))\}_{i\in[d]}\subset\mathbb{R}^{n_{0}}\times\mathbb{R}^{n_{L+1}}$, $d\in\mathbb{N}^{*}$, (i.e., couples of data $x(i)$ and outcomes $y(i):=f(x(i))$) the objective is to produce an estimate of the parameter $\Theta=(\{B_{h}^{(\ell)}\},\{W_{hj}^{(\ell)}\})$, say $\widehat{\Theta}$, in such a way that the output of the neural network is a good estimate of $f$, i.e.,
@@ 2.1. Neural networks and statistical learning
where $\mathcal{T}:=\{(x^{\prime}(j),y^{\prime}(j))\}_{j\in[t]}\subset\mathbb{R}^{n_{0}}\times\mathbb{R}^{n_{L+1}}$, $t\in\mathbb{N}^{*}$ is a test set.
Throughout the paper we encode the inputs (data) in the $n_{0}\times d$ matrix $\mathbb{x}:=(x(1)\ldots x(d))$ and the outcomes (responses) in the $n_{L+1}\times d$ matrix $\mathbb{y}:=(y(1)\ldots y(d))$.
2.2. Neural networks with dependent and possibly heavy-tailed weights, and Bayesian statistical learning
@@
  • For $\ell\in[L+1]$ and $h\in[n_{\ell}]$, $B_{h}^{(\ell)}$ are r.v.’s with $B_{h}^{(\ell)}\sim\mathcal{N}_{1}(0,C_{B}),$ for a constant $C_{B}\geq 0$;
  • For $\ell\in[L+1]$, $h\in[n_{\ell}]$ and $j\in[n_{\ell-1}]$, $W_{hj}^{(\ell)}$ are r.v.’s defined by $W_{hj}^{(\ell)}:=\sqrt{V_{n_{\ell-1},j}^{(\ell-1)}}N_{hj}^{(\ell)},$ where: for $j\in[n_{0}]$, $V_{n_{0},j}^{(0)}:=n_{0}^{-1}$, for $\ell=2,\ldots,L+1$, $\{V_{n_{\ell-1},j}^{(\ell-1)}\}_{j\in[n_{\ell-1}]}$ are non-negative and identically distributed (over $j$) r.v.’s, and for $\ell\in[L+1]$, $h\in[n_{\ell}]$ and $j\in[n_{\ell-1}]$, $N_{hj}^{(\ell)}$ are r.v.’s with $N_{hj}^{(\ell)}\sim\mathcal{N}_{1}(0,C_{W}),$ for a constant $C_{W}>0$;
  • All the r.v.’s $\{B_{h}^{(\ell)},V_{n_{\ell-1},j}^{(\ell-1)},N_{hj}^{(\ell)}\}$ are stochastically independent.
Under these distributional assumptions, one speaks of neural network with dependent and possibly heavy-tailed weights, see the seminal paper [26], where this prior has been introduced. Indeed, note that, for a fixed $\ell\in\{2,\ldots,L+1\}$ and $j\in[n_{\ell-1}]$, the random weights $W_{1j}^{(\ell)},\ldots,W_{n_{\ell}j}^{(\ell)}$ are dependent, and that, if $V_{n_{\ell-1},j}^{(\ell-1)}$ is distributed as the square of a random variable with a heavy tail law with support on $(0,\infty)$, then the distribution of $W_{hj}^{(\ell)}$ is heavy tail. Note also that if $V_{n_{\ell-1},j}^{(\ell-1)}:=\frac{1}{n_{\ell-1}}$, then we recover the usual Gaussian prior. The Bayesian appraoch to the statistical learning problem allows to incorporate in the model the training dataset $\mathcal{D}$ through a likelihood function $\mathcal{L}(\mathcal{D},\Theta)$. Then the posterior knowledge on the parameter is summarized by the posterior probability measure
@@
where $\mathbb{E}_{\rm prior}$ denotes the expectation under the prior probability measure and it is assumed $\mathbb{E}_{\rm prior}\mathcal{L}(\mathcal{D},\Theta)>0.$ If the law of $\Theta$ under the prior has density $\mathrm{p}_{{\rm prior}}(\cdot)$, then the law of $\Theta$ under the posterior has density $\mathrm{p}_{{\rm posterior}}(\theta)\propto\mathcal{L}(\mathcal{D},\theta)\mathrm{p}_{{\rm prior}}(\theta),$ and one estimates the parameter $\Theta$ by maximizing $\mathrm{p}_{{\rm posterior}}(\cdot)$ via an adequate variant of the Stochastic Gradient Descent (maximum a posteriori estimate), see e.g. [16].
3. Preliminaries